# Improving Statistical Significance in Human Evaluation of Automatic Metrics via Soft Pairwise Accuracy

### v1

@@ 4.2 Ablation: Sample Size
Since SPA relies on the pairwise $p$-values between MT systems, it is also natural to ask how SPA behaves when the number of available segments used for evaluating systems is small since it is harder to find statistical differences between systems with a smaller sample size. To answer this question, we calculated 95% confidence intervals for both PA and SPA values of two highly performant metrics—in particular, we considered xCOMET (Guerreiro et al., 2023) and MetricX-23 (Juraska et al., 2023)—on WMT23 using bootstrapping for various numbers of segments, thereby simulating scenarios with less human annotations but a fixed number of MT systems. The results are presented in § 4.2.
@@ 4.3 Ties
As discussed in § 2.2, the quantization noise in PA limits the number of distinct values it can assign to metrics. PA can only take on (${{N\choose 2}}+1$ distinct values. In practice, we find it tends to take on far fewer values. For example, in WMT 22 en-de PA could theoretically take on 92 distinct values, but because the metrics fall in a fairly narrow range of performance (between 62.6% for the worst metric and 81.3% for the best), the 21 metrics have only 11 distinct PA scores, with one 5-way PA tie and several 2- and 3-way PA ties (see Figure 4). Since SPA does not quantize each system comparison, it is able to assign any value to each metric, and is therefore potentially better able to distinguish between metrics.
@@ 4.4 Statistical Significance of Metric Comparisons
Just because a meta-metric produces two distinct scores for two different metrics does not mean that those metrics are statistically significantly different from each other. Ideally, a meta-metric would be able to produce statistically significant comparisons between any two distinct metrics. To explore the ability of PA and SPA to distinguish between metrics in a statistically significant way, we follow PERM-INPUTS Deutsch et al. (2021) to measure the statistical significance of the comparison between each metric. We follow recent shared tasks in greedily computing significance clusters, by starting with the highest scoring metric and assigning rank 1 to all metrics until we encounter the first metric that is statistically different from any previous metric so far. That metric is assigned rank 2, and the process repeats until all metrics have been assigned a rank. We echo the shared task organizers warning that this method can place two metrics that are statistically indistinguishable in different significance clusters (and does so multiple times, in the case of PA). Results are in § 4.4.
@@
On average, SPA increases the number of statistically significant pairwise comparisons by 31% and the number of significance clusters by 40% compared to PA, while producing similar scores for each metric (see Table 1 for results summary, and Figure 4 for a visualization for WMT22 En-De results). This is consistent with our hypothesis that PA is effectively SPA with additive quantization noise.
@@ 5 Historical Context and Related Work
In the WMT 2007-2013 metrics evaluations Callison-Burch et al. (2007, 2008, 2009, 2010, 2011, 2012); Macháček and Bojar (2013) Spearman’s rank correlation coefficient $\rho$ was used for meta-evaluation of metrics. This was motivated by the fact that Spearman’s makes fewer assumptions about the data.
@@
The 2020 metrics task Mathur et al. (2020b) continued to use Pearson’s, but also includes Kendall’s Tau for analysis. Kendall’s Tau is a closer match for the system ranking use case, since it is evaluating if the ordering of a pair of systems is the same as the human ordering. However it is does not take in to account the magnitude difference.
In 2021, the metrics task Freitag et al. (2021) switched to Pairwise accuracy Kocmi et al. (2021), motivated in part by the fact that MT system outliers had an outsized impact on Pearson correlation when it is used to rank MT Metrics Mathur et al. (2020a). The PERM-BOTH hypothesis test of Deutsch et al. (2021) was used to determine significance. 2021 and 2022 Freitag et al. (2022, 2023) follow.
@@ 5.1 Additional Connections to Prior Work
Deutsch et al. (2023) demonstrated that principled tie handling is crucial when comparing MT metrics at the segment (e.g. sentence) level, because some metrics produced quantized scores that commonly create ties. SPA is system level (e.g. sentence level scores averaged over the entire test set), so exact ties are very unlikely. However, SPA can be seen as giving full credit for (statistical) ties, which is similar in spirit.
We show that Quantization is problematic in PA. Quantization in evaluation has proved problematic in other spaces as well—for example, Schaeffer et al. (2024) attributes the widely repeated claim that LLMs have emergent properties to quantization in evaluation.
@@ 6 Conclusions
We introduce a new meta-metric which we denote Soft Pairwise Accuracy, and show that it improves on Pairwise Accuracy in a number of ways, most notably that it is more stable when ablating the number of systems and annotations per system, and that it produces more statistically significant comparisons between metrics. We also discuss how SPA fits into and builds upon the nearly two decade history of meta-evaluation at WMT metric shared tasks.
@@ Acknowledgments
This work would not have been possible without the WMT shared tasks and metrics task releasing their data. We are appreciative of helpful discussions with the WMT metrics task organizers and Rebecca Knowles.
@@
When computing $p$-values, we assume that system-level metric scores are the average of segment-level metrics scores. There is a line of recent work that seeks to incorporate contextual information into automatic metrics. Many such works still produce scores at the segment level (e.g. Vernikos et al., 2022; Hu et al., 2023; Agrawal et al., 2024) but others produce one score per window of a few sentences Raunak et al. (2024) or one score per paragraph Hu et al. (2023). Our method should still be applicable in such cases, but would require permuting windows or paragraphs instead of segments. Additionally, as previously noted, some metrics—notably BLEU Papineni et al. (2002) and chrF Popović (2015)—compute statistics at the segment level and combine them to create document-level scores. Again, permutations would still work but would require some modification. To the best of our knowledge, this issue is not limited to our work - it the same assumption is made in prior work computing statistical significance of metrics, including the WMT shared tasks Freitag et al. (2021, 2022, 2023) and Deutsch et al. (2021).

### v2

@@ 4.2 Ablation: Sample Size
Since SPA relies on the pairwise $p$-values between MT systems, it is also natural to ask how SPA behaves when the number of available segments used for evaluating systems is small since it is harder to find statistical differences between systems with a smaller sample size. To answer this question, we calculate 95% confidence intervals for both PA and SPA values of two highly performant metrics—in particular, we considered xCOMET (Guerreiro et al., 2023) and MetricX-23 (Juraska et al., 2023)—on WMT 2023 using bootstrapping for various numbers of segments, thereby simulating scenarios with less human annotations but a fixed number of MT systems.
@@ 4.3 Ties
[…]
@@ 4.4 Statistical Significance of Metric Comparisons
We hypothesize that the binarization in PA is essentially acting as additive random noise on top of the underlying SPA meta-metric. If this is true (and the magnitude of the noise does not dominate the underlying signal), we would expect SPA to produce a similar metric ranking to PA, but with increased statistical significance. To test this, we compute statistical significance of the comparisons between each metric using the PERM-INPUTS Deutsch et al. (2021) method. We follow recent shared tasks in greedily computing significance clusters, by starting with the highest scoring metric and assigning rank 1 to all metrics until we encounter the first metric that is statistically significantly different from any previous metric so far. That metric is assigned rank 2, and the process repeats until all metrics have been assigned a rank. We echo the shared task organizers’ warning that this method can place two metrics that are statistically indistinguishable in different significance clusters (and in the case of PA, we observe this multiple times).
@@
[…]
@@ 5 Historical Context and Related Work
In the WMT 2007-2013 metrics evaluations Callison-Burch et al. (2007, 2008, 2009, 2010, 2011, 2012); Macháček and Bojar (2013) Spearman’s rank correlation coefficient $\rho$ was used for meta-evaluation of metrics. This was motivated by the fact that Spearman’s makes fewer assumptions about the data than the Pearson correlation coefficient.
@@
The 2020 metrics task Mathur et al. (2020b) continued to use Pearson’s, but also includes Kendall’s Tau for analysis. Kendall’s Tau is a closer match for the system ranking use case, since it is evaluating whether the ordering of a pair of systems is the same as the human ordering. However, it does not take into account the magnitude difference.
In 2021, the metrics task Freitag et al. (2021) adopted pairwise accuracy Kocmi et al. (2021), motivated in part by the fact that MT system outliers had an outsized impact on Pearson correlation when it is used to rank MT Metrics Mathur et al. (2020a). Pairwise accuracy produces the same system-level ranking as Kendall’s Tau, as they are equivalent modulo a linear scaling and shifting (see § 5.1). The PERM-BOTH hypothesis test of Deutsch et al. (2021) was used to determine significance. 2021 and 2022 Freitag et al. (2022, 2023) follow.
@@
5.1 Relationship to Kendall’s Tau
Our work builds on pairwise accuracy, typically attributed to Kocmi et al. (2021). Pairwise accuracy is equivalent to the the widely used Kendall rank correlation coefficient Kendall (1938), modulo a linear scaling and shifting. Kocmi et al. (2021) present pairwise accuracy as simply “accuracy” and make no mention of its relation to Kendall, which was already in use for MT meta-evaluation Mathur et al. (2020b). The term “pairwise accuracy” appears to have been coined by Freitag et al. (2021) to distinguish it from other types of accuracy.
Kendall’s Tau is defined in terms of concordance (equivalent to our previously defined $a_{ij}^{m}$) and discordance $\,d_{ij}^{m}$, defined to be 1 when the metric and human judgments disagree and 0 otherwise:
$\tau$ =($={{N\choose 2}}^{-1}\,\,\sum_{i=0}^{N-1}\sum_{j=i+1}^{N-1}\,(a_{ij}^{m}-d_{ij}^{m})$ (4)
Any system pair which is not concordant is discordant, [^7] We ignore tie handling, as ties are extremely unlikely in system-level evaluation. Ties in segment-level evaluation are an entirely different matter Deutsch et al. (2023a). and thus $d_{ij}^{m}={1}-a_{ij}^{m}$. Given this and the definition of PA from Equation 2, we have:
$\begin{split}\tau&={{N\choose 2}}^{-1}\,\,\sum_{i=0}^{N-1}\sum_{j=i+1}^{N-1}\,a_{ij}^{m}-({1}-a_{ij}^{m})\\ &=2\Bigg{(}{{N\choose 2}}^{-1}\,\,\sum_{i=0}^{N-1}\sum_{j=i+1}^{N-1}\,a_{ij}^{m}\Bigg{)}-1\\ &=2\;{PA}-1\end{split}$ (5)
@@ 5.2 Additional Connections to Prior Work
Deutsch et al. (2023a) demonstrated that principled tie handling is crucial when comparing MT metrics at the segment level, because some metrics produce quantized scores that often result in ties. SPA is system level (i.e. sentence level scores averaged over the entire test set), so exact ties are very unlikely. However, SPA can be seen as giving full credit for (statistical) ties, which is similar in spirit.
We show that quantization (specifically binarization) is problematic in PA. Quantization in evaluation has proved problematic in other spaces as well—for example, Schaeffer et al. (2024) attributes the widely repeated claim that LLMs have emergent properties to quantization in evaluation.
@@ 6 Conclusions
We introduce a new meta-metric which we denote soft pairwise accuracy, and show that it improves on pairwise accuracy in a number of ways, most notably that it is more stable than pairwise accuracy when ablating the number of systems and annotations per system, it fixes an issue of metric ties observed in pairwise accuracy, and it produces more statistically significant comparisons between metrics than pairwise accuracy. We also discuss how soft pairwise accuracy fits into and builds upon the nearly two decade history of meta-evaluation at the WMT Metric Shared Tasks.
@@ Acknowledgments
This work would not have been possible without the WMT shared tasks and metrics tasks releasing their data. We are appreciative of helpful discussions with the WMT metrics task organizers as well as Sweta Agrawal and Rebecca Knowles.
@@
When computing $p$-values, we assume that system-level metric scores are the average of segment-level metrics scores. There is a line of recent work that seeks to incorporate contextual information into automatic metrics. Many such works still produce scores at the segment level (e.g. Vernikos et al., 2022; Hu et al., 2023; Agrawal et al., 2024) but others produce one score per window of a few sentences Raunak et al. (2024) or one score per paragraph Deutsch et al. (2023b). Our method should still be applicable in such cases, but would require permuting windows or paragraphs instead of segments. Additionally, as previously noted, some metrics—notably BLEU Papineni et al. (2002) and chrF Popović (2015)—compute statistics at the segment level and combine them to create document-level scores. Again, permutations would still work but would require some modification. To the best of our knowledge, this issue is not limited to our work—the same assumption is made in prior work computing statistical significance of metrics, including the WMT shared tasks Freitag et al. (2021, 2022, 2023) and Deutsch et al. (2021).